OpenStudy (gavin39):

Write an equation of the line. Horizontal: through (0,-4) Equation of the line is?

OpenStudy (gavin39):

@Owlcoffee

OpenStudy (gavin39):

@MarcoReus11

OpenStudy (gavin39):

I think it's y=-4

OpenStudy (sshayer):

horizontal line has slope=0 eq . of line is y-y1=m(x-x1) y-(-4)=0(x-0) y+4=0 y=-4 so you are correct.

OpenStudy (gavin39):

@sshayer Thank you. Could you help me with another I was having trouble with?